Consider the following equation "I=\displaystyle{\int_{0}^{1}\sin{(x^2)}dx}"

Find the integral  "\displaystyle{\int_{-1}^{1}\sin{(x^2)}dx}" in terms of "I"


1
Expert's answer
2021-09-23T00:37:12-0400

"\sin x = \sum \dfrac{(-1)^n x^{2n+1}}{(2n+1)!}"


"\text{then} \; \; \sin (x^2) = \sum \dfrac{(-1)^n x^{4n+2}}{(2n+1)!}"


"\int \sin (x^2) = \int \sum \dfrac{(-1)^n x^{4n+2}}{(2n+1)!} = \sum \dfrac{(-1)^n x^{4n+3}}{(2n+1)!(4n+3)} + K"


"\therefore\int_0^1\sin x^2~dx"


"=\left[\sum\limits_{n=0}^\infty\dfrac{(-1)^nx^{4n+3}}{(2n+1)!(4n+3)}\right]_0^1"


"=\sum\limits_{n=0}^\infty\dfrac{(-1)^n}{(2n+1)!(4n+3)}=I"


"\int_{-1}^1\sin x^2~dx"


"=\left[\sum\limits_{n=0}^\infty\dfrac{(-1)^nx^{4n+3}}{(2n+1)!(4n+3)}\right]_{-1}^1"


"=\sum\limits_{n=0}^\infty(2\dfrac{(-1)^n}{(2n+1)!(4n+3)})= 2I"